OpenStudy (anonymous):

At 2:45 p.m you are 112 miles from Dallas. You want to be in Dallas at 4:30 p.m what is the average speed you must travel to be on time ?

OpenStudy (phi):

can you figure out how much time between 2:45 and 4:30 ?

OpenStudy (anonymous):

v= d /t v= 112 / (4:30 -2:45) v=112 / 1.75 h v=64mile/h